a person with normal near point 25 cm using a compound microscope with objective of focal length 8 mm and an eye piece of focal length 2.5 cm can bring an object placed at 9 mm from the objective in sharp focus.the separation between two lenses and magnification respectively are
1]9.47cm,96
2]3.36cm,44
3]6cm,22
4]7.49cm,11 

Grade:8

1 Answers

Raveesh
7 years ago
For sharp focus, image should be formed at infinity.
For image to be at infinity, The image formed by the objective should be at the focus of the eyepiece.
WKT,
\frac{1}{f}                             
Magnification(m)= vo/uo(D/fe)
m=2.5/0.9(25/2.5)
m=2.5(10)/0.9
m=250/9